aliyah32:

You have $1,998.00 on a credit card with a 14.5% APR. You miss your minimum payment the first month and there is a late fee of $37.00. How much is your balance at the beginning of the second month? $2,083.29 $2,046.29 $2,022.14 $2,059.14

aliyah32:

??

aliyah32:

??

aliyah32:

??

aliyah32:

bruh

aliyah32:

you gonna answer or not?

Convert:

$1,998*0.145 (converting to a decimal, moving 2 places to the left)=$289.71 $1,998+$289.71=$2287.71 $2287.71+$37=$2324.71 The answer would be $2324.71 but I do not see an answer with that

aliyah32:

that is not right

aliyah32:

there is no option for answer

Convert:

oh your right its $1998. x (.145/12) + 37 = 2,059.14 so the last answer

Convert:

sorry

hiiiihiiiiiiiiii:

iiugyftrdzawsxedcrfvtgbyhujomkjnihbugvfcdxszAQ@wsedftgyhujjjjhgfdsazxcvbhhgfdswe45678

MxxnLight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert sorry \(\color{#0cbb34}{\text{End of Quote}}\) It's okay at least you tried.

MxxnLight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hiiiihiiiiiiiiii iiugyftrdzawsxedcrfvtgbyhujomkjnihbugvfcdxszAQ@wsedftgyhujjjjhgfdsazxcvbhhgfdswe45678 \(\color{#0cbb34}{\text{End of Quote}}\)

aliyah32:

thanksss, i appreciate it.

Convert:

No problem, glad to help! :)

supie:

Amount on the Credit card was \[$1,998.00\] APR = \[14.5 %\] Monthly APR = \[14.5/12=1.2083\] Amount that is due after a month that wasn't paid \[=1.2083 \ of \ $ 1998.00\] \[1.2083(1998.00)\] \[=24.14\] Late fees = \[37.00\] Amount that needs to be paid at the beginning of Second month = Actual money in credit card + Interest + Late fees \[= 1998 + 24.14 + 37\] \[= 2035 +24.14\] \[= $2059.14\]